OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 1 4 5 25

OpenStudy (katrinakaif):

Distribute 3 first to (x-1)

OpenStudy (anonymous):

i did is it 25 ?

OpenStudy (anonymous):

thank you brinethery and katrina i got more need help with so ima post them look out for them

OpenStudy (anonymous):

i did thatbut is it 25 ?

OpenStudy (katrinakaif):

Your first step is to always distribute: In this case, you have two terms to distribute: 3 to x + 1 -2 to x -1 It basically means: 3(x + 1) ---> (3 x 'x') + (3 x 1) and -2 (x - 1) ---> (-2 x 'x') - (-2) x (-1)

OpenStudy (anonymous):

3x+3 = -2x +2 +6 Can you solve this?

OpenStudy (anonymous):

no

OpenStudy (katrinakaif):

Your goal for linear functions is to get all the constants to one side, and all variables to the other. Try it

OpenStudy (anonymous):

im not asking yall to show me how to do it im just trying to check my answers

OpenStudy (anonymous):

You told me that you didn't know how to solve this. 3x+3 = -2x +2 +6

OpenStudy (katrinakaif):

You got 25 yes?

OpenStudy (katrinakaif):

Patience is a virtue, is it not?

OpenStudy (anonymous):

yes i got 25

OpenStudy (anonymous):

@VRFGVD I know you don't like to hear this, but you need to learn how to solve for x before you start learning how to distribute stuff. It's all a stepping stone thing.

OpenStudy (katrinakaif):

That is inaccurate Vrfgvd. Lets try again: Can you start us off with the first step?

OpenStudy (anonymous):

3x+3 = -2x +8 What do you do first? I would subtract 3 from both sides.

OpenStudy (anonymous):

And then add 2x to both sides.

OpenStudy (anonymous):

so 10x

OpenStudy (katrinakaif):

Why? Because remember, in a linear function, variables go on one side of the equal sign, and constants on the other.

OpenStudy (anonymous):

divide that by 2 then it leaves 5

OpenStudy (anonymous):

3x +3 -3 = -2x +8 -3 3x = -2x +5 3x +2x = -2x +2x +5 5x = 5

OpenStudy (anonymous):

i need help with this one to please Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (katrinakaif):

Would you like to use twiddla to have a better understanding?

OpenStudy (anonymous):

@VRFGVD what grade level is this for?

OpenStudy (anonymous):